Home‎ > ‎Reports‎ > ‎Crew103‎ > ‎

day09


Crew 103 Log Book for April 17, 2011

Photos of the day





Commander's Check-In Report

Anne-Sophie Kalbfleisch Reporting

Crew Physical Status: Everybody is fine

Time Departed/Returned from EVA:

EVA 1 : 10h-10h30

EVA 2 : 15h30-17h45

Brief Narrative of Field Mission Results:

During the first EVA, G. Van Roy & J. Offergeld went outside to re-position the antenna, a higher distance from the internet antenna, in order to have less noise on the received signal. They were accompanied by the max rover.

During the afternoon, I took some measurements of our experiment with the algae. Three other crew-members went on EVA in order to collect more soil samples for the permeability analysis.

EVA Data/Interpretations:

We didn't notice less noise on the radio-telescope's signal.

Concerning the algae, it's difficult to see any evolution after a short time.

The soil analysis is going on.

Engineering/Hab Maintenance:

nothing to mention

Plans for Tomorrow:

We will continue the measurements of our algae.

We would like to repair the greywater system but first we will ask Joshua if it's possible. It seemed that it didn't work anymore and Don Lusko switched the circuit off.

Inventory:

nothing to mention

Support Requested:

Some help from Joshua for a possible repair to the greywater filtering system. We would like to know more about that (why precisely isn't it

working anymore?) and what we could possibly do.

Miscellaneous:

Nothing to mention


Commander's Report

Anne-Sophie Kalbfleisch Reporting

Today it was Sunday so we decided to have a more calm day and rest a little. During the morning G. Van Roy & J. Offergeld went outside to

move the radio-telescope antenna to another place because we were afraid that the internet antenna influenced the signal by producing some noise. They went outside with the max rover. But this hypothesis proved false because we found that moving the antenna didn't reduce the noise.

We did also one hour psychological tests for the psychologist who is following us from Belgium. When we get the results of her study we will send it back to Mission Support but that will be after our stay at MDRS.

In the afternoon, G. Van Roy, W. de Groot & J. Offergeld went outside with the ATV's for a second EVA in order to continue our permeability

study by collecting and analyzing soil samples.

I also took some measurements for our algae experience. We are very happy because the optical density increased, which means that our little

algae are growing.

Plans for Tomorrow:

We will continue the measurements of our algae.

We would like to repair the greywater system but first we will ask Joshua if it's possible. It seemed that it didn't work anymore and Don

Lusko switched the circuit off.

Inventory:

N/A

Support Requested:

Some help from Joshua for a possible repair to the greywater filtering system. We would like to know more about that (why precisely isn't it

working anymore?) and what we could possibly do.


Engineering Report

Thibault Windels Reporting

Generator/Electricity:

Honey (Diesel Generator) (Used? Yes/No): yes

Wendy (Gas Generator) (Used? Yes/No): no

Battery Charge Level: 10,3 volts

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off): generator remained on


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0,5

Propane Tank (Full/0.75/0.5/0.25/Empty): 0,39

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 6 gallons approximately

Notes/Comments: /


Oil:

Kitty Oil Quantity (# of Quarts in storage): the previous commander told us not to check the oil quantity of Kitty because we have to turn the generator off to do so!

Honey Oil Quantity (# of Quarts in storage): /

ATV Oil Quantity (# of Quarts in storage): full

Car Oil Quantity (# of Quarts in storage): /


Water Status:

Outside Potable Water Tank Level (inches from bottom): 14,5

Trailer Potable Water Tank Level (inches from bottom): 12,5

Hab Potable Water Tank Level (inches from bottom): 5

Potable Water Meter Reading: 45240,3

In to GreenHab Meter Reading: /

Notes/Comments: /


GreenHab:

Condition of plants in Tank 1: /

Condition of plants in Tank 2: /

Condition of plants in Tank 3: /

Greenhab Notes/Comments: One of the two ventilators is on, I don't find the electric source for the other. The thermostat is at 50°F and the temperature inside is 86°F


Observatory:

Telescope Used? (Yes/No): no

Observatory Notes/Comments: there's no telescope


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no): no

Tire Status: ok

Odometer at end of day: 124250,5

Notes/Comments on Hab Car: /


ATV 1 (Opportunity):

Used: yes

Oil Checked: no

Fuel consumed : 1/4 tank approximately

Tire Status: ok

Comments: /


ATV 2 (Spirit):

Used: yes

Oil Checked: no

Fuel Consumed: 1/4 tank approximately

Tire Status: ok

Comments: /


ATV 3 (Viking I):

Used: yes

Oil Checked: no

Fuel Consumed: 1/4 tank approximately

Tire Status: ok

Comments: /


ATV 4 (Viking II):

Used: no

Oil Checked: no

Fuel Consumed: 0

Tire Status:

Comments: does not work


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 60

Actual temperature upstairs (Farenheit): 79

Thermostat setting downstairs (Farenheit): /

Actual temperature downstairs (Farenheit): 75


Computers/Networking Infrastructure:

Notes/comments on internet/computers: Internet is still running well


General Engineering Notes/Comments: Could we possibly repair the grey water system?


Science Report (Geology)

Wouter de Groot Reporting

Today during the second EVA we made new measures of the soil permeability. We followed the same procedure that Grégory described two days ago.

Location 1 :

Coordinates : 12S 0519861 / 4251222

Plants : very weak vegetation, herbaceous

Soil condition : dry

Soil type : sand more pebbles, near dried river bed

Soil temperature : 33°C

Permeability measure : 2 inch / 1 min => 304,8 cm / hour => very high

Location 2 :

Coordinates : 12S 0520165 / 4248950

Plants : Moderate

Soil condition : dry

Soil type : sand, between two dried river beds

Soil temperature : 32°C

Permeability measure : 2,6 inch / 2 min => 198,12 cm / hour => very high

Location 3 :

Coordinates : 12S 0520182 / 4248937

Plants : Some herbaceous but almost nothing

Soil condition : dry

Soil type : red with a crust above and some pebbles

Soil temperature : 32°C

Permeability measure : 0,4 inch / 6 min => 10,16 cm / hour => moderate

Location 4 :

Coordinates : 12S 0515196 / 4251246

Plants : Almost nothing

Soil condition : dry

Soil type : compact sand

Soil temperature : 26°C

Permeability measure : 1 inch / 10 min => 15,24 cm / hour => high



EVA 6 Report

Gregory Van Roy Reporting

Time: 10 - 10.30 am

Crew Members: Jonas Offergeld and Gregory Van Roy

Site Location: Front of the station

Transit Mode: by foot

Objectives: Move the antenna of the radio-telescope

Results: The antenna and the spectrum analysis work better than yesterday

Lessons Learned: None


EVA 7 Report

Wouter de Groot Reporting

Time: 3.30 pm - 5.45 pm

Crew Members:  Jonas Offergeld, Gregory Van Roy and Wouter de Groot

Site Location: 12S 0519861 / 4251222

                      12S 0520165 / 4248950

                      12S 0520182 / 4248937

                      12S 0515796 / 4251246


Transit Mode: by ATV

Objectives: make new soil permeability measures

Results: See the Science report

Lessons Learned: None


Chef's Report

Gregory Van Roy Reporting

1a) Today was

A non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

Nothing special today

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.

Bear Creek  Cheddar Broccoli Soup,Thrive Rice/Dijon Chicken, m&m's cookies and Strawberry drink mix

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

None

4)  Recipes for anything more complicated than following package directions exactly, or re-hydrating.

None

5)  Lessons learned  (if any)

None

6) Comments/questions for Kim and Jean, the food study investigators

We can use more Quaker instant oatmeal and more sweet things because we used less than the half of the reserve in a week.